OpenStudy (anonymous):

wat is -5/12 + 2/5 pleez show work

OpenStudy (anonymous):

-5/12 + 2/5 here we have to first find the LCM of 12 and 5 that is 60 (-5(5) + 2(12) ) /60 = (-25 + 24)/60 = -1/60 hope u understoood

OpenStudy (anonymous):

multiply the first fraction by 5/5 and the second by 12/12. -5/12 * 5/5 + 2/5 * 12/12 -25/60 + 24/60 = -1/60

OpenStudy (anonymous):

did u understand it Andreacurry